# Freezing-in the Axiverse

### v1

@@ Appendix A General Considerations of Axiverse EFT
The Lagrangian containing $\cal N$} axions $\theta_{i=1,\cdots,{\cal N}}$ is given by (see, e.g.Mehta:2021pwf; Gendler:2023kjt),
@@
where the indices $A$, II, $\{i,j\}$ run over the SM operators, instantons generating axion potentials, and the different axions, respectively. ${\cal O}^{(5)}_{A}$ and ${\cal O}^{(6)}_{A}$ are the various SM operators that determine dimension-5 and dimension-6 axion-SM couplings, respectively. For example, for the gluon coupling ${\cal O}^{(5)}_{G}=\alpha_{s}G\tilde{G}/(8\pi)$, for the fermion coupling, ${\cal O}^{(5)}_{\Psi}=\partial_{\mu}(\bar{\Psi}\gamma^{\mu}\Psi)$ (where we have suppressed the flavor indices), and for the dimension-6 coupling to the Higgs ${\cal O}^{(6)}_{A,\mu\nu}=|H|^{2}g_{\mu\nu}$. For each SM operator $A$, $\tilde{c}_{A}$ and $\tilde{d}_{A}$ are $\cal N$}-vector and ${\cal N}\times{\cal N}$ matrix, respectively. Both the instanton charge matrix $\bar{Q}_{Ij}$ and anomalous gauge boson couplings consist of integers such that ${\cal L}_{\rm Axiverse}$ enjoys discrete shift symmetries $\theta_{i}\rightarrow\theta_{i}+2\pi$. For concreteness, we have modeled the periodicity of the axion potential by a cosine. This axion potential can come from some UV physics, such as confining dark gauge groups or charged particles in extra dimensional theories Arkani-Hamed:2003xts. Furthermore, in a given model it could be some other periodic function $f(\bar{Q}_{Ij}\theta_{j}+\delta_{I})$, and the next discussion does not rely on the precise nature of this function. We will focus on the case where the number of instantons is larger than the number of axions $\cal N$}, as seen in some string theory constructions Mehta:2021pwf.
It is possible to write ${\cal L}_{\rm Axiverse}$ in a more convenient form, under the assumption that the sizes of the instanton-generated potentials are hierarchical, $\epsilon_{I}\equiv\Lambda_{I+1}^{4}/\Lambda_{I}^{4}\ll 1$ Gendler:2023kjt. With this assumption, we can select $\cal N$} leading instantons with a charge matrix $Q_{ab}$, obtained by removing any row from $\bar{Q}_{Ia}$ that is a linear combination of the rows above it. We can also use $2\pi$-shifts of $\theta_{i}$ to remove a certain number of arbitrary phases $\delta_{I}$. With this procedure, the axion potential from ${\cal L}_{\rm Axiverse}$ can be written as Gendler:2023kjt,

### v2

@@ Appendix A General Considerations of Axiverse EFT
The Lagrangian containing $\cal N$} axions $\theta_{i=1,\cdots,{\cal N}}$ is given by (see, e.g.Mehta et al. (2021); Gendler et al. (2024)),
@@
where the indices $A$, II, $\{i,j\}$ run over the SM operators, instantons generating axion potentials, and the different axions, respectively. ${\cal O}^{(5)}_{A}$ and ${\cal O}^{(6)}_{A}$ are the various SM operators that determine dimension-5 and dimension-6 axion-SM couplings, respectively. For example, for the gluon coupling ${\cal O}^{(5)}_{G}=\alpha_{s}G\tilde{G}/(8\pi)$, for the fermion coupling, ${\cal O}^{(5)}_{\Psi}=\partial_{\mu}(\bar{\Psi}\gamma^{\mu}\Psi)$ (where we have suppressed the flavor indices), and for the dimension-6 coupling to the Higgs ${\cal O}^{(6)}_{A,\mu\nu}=|H|^{2}g_{\mu\nu}$. For each SM operator $A$, $\tilde{c}_{A}$ and $\tilde{d}_{A}$ are $\cal N$}-vector and ${\cal N}\times{\cal N}$ matrix, respectively. Both the instanton charge matrix $\bar{Q}_{Ij}$ and anomalous gauge boson couplings consist of integers such that ${\cal L}_{\rm Axiverse}$ enjoys discrete shift symmetries $\theta_{i}\rightarrow\theta_{i}+2\pi$. For concreteness, we have modeled the periodicity of the axion potential by a cosine. This axion potential can come from some UV physics, such as confining dark gauge groups or charged particles in extra dimensional theories Arkani-Hamed et al. (2003). Furthermore, in a given model it could be some other periodic function $f(\bar{Q}_{Ij}\theta_{j}+\delta_{I})$, and the next discussion does not rely on the precise nature of this function. We will focus on the case where the number of instantons is larger than the number of axions $\cal N$}, as seen in some string theory constructions Mehta et al. (2021).
It is possible to write ${\cal L}_{\rm Axiverse}$ in a more convenient form, under the assumption that the sizes of the instanton-generated potentials are hierarchical, $\epsilon_{I}\equiv\Lambda_{I+1}^{4}/\Lambda_{I}^{4}\ll 1$ Gendler et al. (2024). With this assumption, we can select $\cal N$} leading instantons with a charge matrix $Q_{ab}$, obtained by removing any row from $\bar{Q}_{Ia}$ that is a linear combination of the rows above it. We can also use $2\pi$-shifts of $\theta_{i}$ to remove a certain number of arbitrary phases $\delta_{I}$. With this procedure, the axion potential from ${\cal L}_{\rm Axiverse}$ can be written as Gendler et al. (2024),